# PaddleOCR-VL: Boosting Multilingual Document Parsing via a 0.9B Ultra-Compact Vision-Language Model

### v3

@@ 1 Introduction
Documents serve as core information carriers, with their complexity and volume growing at an exponential rate, making document parsing an indispensable key technology. The primary goal of document parsing [1, 2, 3, 4] is to enable deep structural and semantic understanding of a document’s layout. Specifically, it involves recognizing distinct text blocks and columns, distinguishing formulas, tables, charts, and images, determining the correct reading order, and detecting key elements (e.g., footnotes and image captions); these capabilities collectively lay a solid foundation for efficient information retrieval and data management. Furthermore, advanced document parsing enables large language models (LLMs) [5, 6, 7], especially when combined with Retrieval-Augmented Generation (RAG) [8], to access high-quality knowledge and enhance their practical applications.
The inherent complexity of modern documents presents unique challenges: they often combine dense text, complex tables or chart, mathematical expressions, multiple languages and handwritten texts, with diverse layout structures. Recent research [1, 9, 10, 11, 12] in the field of document parsing primarily following two technological approaches. The first approach [9, 10] employs pipeline methodologies based on specialized, modular expert models. Although these methods offer strong performance, they are increasingly hindered by integration complexity, cumulative error propagation, and inherent limitations when handling highly complex documents. Secondly, end-to-end approaches [12, 13, 14] leveraging multimodal models aim to simplify the workflow and enable joint optimization. However, these methods often struggle with correct text order and can even generate hallucinations when faced with lengthy or complex layouts, while also incurring substantial computational overhead for long sequence outputs, thereby restricting their practical deployment.
@@
Firstly, PaddleOCR-VL performs layout detection and reading order prediction to obtain the positional coordinates and reading order of elements (text blocks, tables, formulas, and charts). Compared to multimodal methods that rely on grounding and sequence output (e.g., MinerU2.5 [2], Dolphin [3]), our method offers faster inference speeds, lower training costs, and easier extensibility for new layout categories. Subsequently, the elements are segmented based on their positions and fed into PaddleOCR-VL-0.9B for recognition. This vision-language model is specifically designed for resource-efficient inference and excels at element recognition within document parsing. By integrating a NaViT-style [15] dynamic high-resolution visual encoder with the lightweight ERNIE-4.5-0.3B [5] language model, we have significantly enhanced the model’s dense text recognition capabilities and decoding efficiency.
@@
Comprehensive benchmarking on the public benchmarks, including OmniDocBench v1.0, v1.5 [16] and olmOCR-Bench [12], and in-house ones demonstrate that PaddleOCR-VL achieves SOTA performance in document parsing task, significantly outperforming existing pipeline-based solutions and exhibiting strong competitiveness against leading vision-language models (VLMs). Moreover, PaddleOCR-VL is optimized for efficiency, delivering substantially lower latency and higher throughput than competing approaches.
@@ 2.1.1 Layout Analysis
Specifically, we decouple the layout analysis process by introducing an independent model, PP-DocLayoutV2, dedicated solely to this task. PP-DocLayoutV2 consists of an object detection model (RT-DETR [17]) for elements localization and classification, as well as a lightweight pointer network [18] with six transformer layers to accurately predict the reading order of layout elements.
@@
Architecturally, PP‑DocLayoutV2 is composed of two sequentially connected networks, as shown in Figure 3. The first is an RT‑DETR-based [17] detection model that performs layout element detection and classification. The detected bounding boxes and class labels are then passed to a subsequent pointer network, which is responsible for ordering these layout elements.
Specifically, we first apply per-class thresholds to select foreground proposals for the ordering network. The selected proposals are embedded using absolute 2D positional encodings and class label embeddings. Additionally, the encoder attention incorporates a geometric bias mechanism from Relation‑DETR [18] to explicitly model pairwise geometric relationships among elements. The pairwise relation head linearly projects element representations into query and key vectors, then computes bilinear similarities to produce pairwise logits, resulting in an $N\times N$ matrix that represents the relative order between each pair of elements. Finally, a deterministic win‑accumulation decoding algorithm recovers a topologically consistent reading order for the detected layout elements.
In comparison to other specialized models, such as LayoutReader [19], our model achieves higher performance with fewer parameters by efficiently extending RT-DETR [17] with a pointer network.
@@ 2.1.2 Element-level Recognition
We adopted an architectural style inspired by LLaVA [20], integrating a pre-trained vision encoder with a dynamic resolution preprocessor, a randomly initialized 2-layer MLP projector, and a pre-trained large language model. Our architecture achieves a balance the scale of vision and language models to optimize performance in multi-elements recognition tasks.
Compared to earlier document parsing models based on fixed-resolution or tiling-based approaches [4, 14, 21], our approach utilizes native dynamic high-resolution preprocessing. For the vision encoder, we employed a NaViT-style [15] encoder initialized from Keye-VL’s [22] vision model, which support native-resolution inputs. This design enables the vision-language model to handle images of arbitrary resolution without distortion, yielding fewer hallucinations and stronger performance on text-intensive tasks.
The projector is a randomly initialized 2-layer MLP with GELU [23] activation, incorporating a merge size of 2 to efficiently bridge visual features from the encoder to the language model’s embedding space.
In auto-regressive language models, the entire sequence is generated by predicting one token at a time. This approach means that the size of the decoder is directly linked to the overall inference latency, so a smaller model will decode faster. With this in mind, we use the ERNIE-4.5-0.3B [5] model, an open-source language model that balances a relatively small number of parameters with strong inference efficiency. In our implementation, we further enhance positional representation by incorporating a 3D-RoPE[24].
The combination of NaViT [15] with ERNIE-4.5-0.3B [5] has led to significant performance improvements in documents parsing, achieving minimal memory usage and faster inference speed.
@@ 2.2.1 Layout Analysis
We employ the PP-DocLayoutV2 model to perform layout element localization, classification, and reading order prediction. PP-DocLayoutV2 extends RT-DETR [17] by incorporating an additional pointer network [18], which is responsible for predicting the reading order of detected elements. The training process adopts a two-stage strategy: we first train the core RT-DETR [17] model for layout detection and classification. Afterward, we freeze its parameters and independently train the pointer network for reading order prediction.
For the first stage, we follow the training strategy of RT-DETR [17]. Specifically, we initialize the model with PP-DocLayout_Plus-L [25] pretrained weights and train it for 100 epochs on our self-constructed dataset comprising over 20,000 high-quality samples.
For the second stage, specifically, the model outputs a matrix representing the pairwise ordering relationships between any two elements, and the Generalized Cross Entropy Loss [26] is computed with respect to the ground truth labels, as this loss function demonstrates increased robustness in scenarios where pre-annotated data are mixed into the dataset. We utilize a constant learning rate 2e-4 and the AdamW optimizer to train 200 epochs.
@@ 2.2.2 Element-level Recognition
As described in Section 2.1.2, PaddleOCR-VL-0.9B consists of three modules: a vision encoder, a projector, and a language model. We adopt a post-adaptation strategy using pre-trained models. Specifically, the vision model is initialized with Keye-VL’s weights, and the language model is initialized with ERNIE-4.5-0.3B’s weights. The model is trained based on the ERNIEKit [27] repository and the training methodology for our VLM is divided into two stages, as outlined in Table 1.
@@
    Table Recognition: The model learns to parse tabular structures within documents. This involves accurately extracting cell contents, identifying rows and columns, and recognize the logical relationships between different table elements, ultimately generating structured representations based on OTSL [28] format.
@@ 3.1 Data Curation
    Open Source Dataset: As the foundation of our dataset, we systematically aggregated and curated a wide array of established public datasets. For textual content, we sourced data from the canonical dataset CASIA-HWDB [29]. Our mathematical expression data is derived from UniMER-1M [30] and MathWriting [31]. To ensure comprehensive coverage of data visualizations, we incorporated a rich spectrum of chart and graph datasets, including ChartQA [32], PlotQA [33], Chart2Text [34], DVQA [35], Unichart [36], Beagle [37], ChartINFO [38], visText [39], and ExcelChart [40]. Each of these sources underwent an initial filtering and cleaning protocol to rectify or discard noisy and low-quality annotations.
@@ 3.2 Automatic Data Annotation
After acquiring the raw data, we utilize an automatic data annotations process for large-scale labeling. Initially, we employ the expert model, PP-StructureV3, to conduct preliminary processing on the data, generating pseudo labels that may contain some inaccuracies. Subsequently, through prompt engineering, we create prompts that include the original images and their associated pseudo labels, which are then submitted to more advanced multimodal large language models, ERNIE-4.5-VL [5] and Qwen2.5VL [24]. These sophisticated models refine and enhance the initial results by analyzing the image content, resulting in improved labels. Finally, to ensure the quality of the labels, the system performs a hallucination filtering step, which eliminates any potentially incorrect content generated by the large models, thereby producing reliable and high-quality labels.
@@ 3.3 Hard Cases Mining
To overcome performance bottlenecks in specific complex scenarios, we propose a hard case mining process for targeted performance improvement. We firstly develop a eval engine for various types. We created substantial evaluation data with precisely labeled data obtained through manual annotation. Theses evaluation datasets are categorized into several types: text data includes 23 categories such as Chinese, English, printed, handwritten, Japanese, Latin, and emojis; table data includes 20 categories such as limited tables, unlimited tables, handwritten tables, checklists, invoices, and rotated tables; formula data includes 4 categories such as Chinese and English formulas, handwritten and printed, simple, and complex; chart data includes 11 categories such as Chinese and English charts, line charts, and bar charts, sourced from diverse origins to cover different document. By inference on this evaluation set and using corresponding professional metrics (e.g., EditDist for Text, TEDS [41] for Tables, RMS-F1 [42] for Charts, and BLEU [43] for Formulas), we can accurately identify hard cases where the model performs poorly. Finally, for these identified weaknesses, the system utilizes a rich set of resources (such as Font Library, CSS Library, Corpus) and rendering tools (like XeLaTeX and web browsers) to synthetically generate a large volume of new, high-quality hard cases.
@@ 4.1 Page-level Evaluation
To comprehensively evaluate the document parsing capabilities, we conducted extensive experiments on the OmniDocBench v1.5 [2] benchmark. It is an expansion of version v1.0, adding 374 new documents for a total of 1,355 document pages. It features a more balanced distribution of data in both Chinese and English, as well as a richer inclusion of formulas and other elements. The evaluation method has been updated, with formulas assessed using the CDM method. The overall metric is a weighted combination of the metrics for text, formulas, and tables.
Table 2 demonstrate that PaddleOCR-VL achieves SOTA performance, outperforming existing pipeline tools, general VLMs, and other specialized document parsing models across all key metrics. Specifically, our model achieves a top-ranking overall score of 92.86, surpassing the next best model, MinerU2.5-1.2B (90.67). Moreover, our model establishes new SOTA results in the sub-tasks, including the lowest Text-Edit distance [44] of 0.035, the highest Formula-CDM score of 91.22, the leading scores of 90.89 and 94.76 in Table-TEDS and Table-TEDS-S, and the best readering ordering scores of 0.043, respectively. These results underscore its superior accuracy in text recognition, formula recognition, and complex table structure analysis.
@@
Pipeline Tools Marker-1.8.2 [45] - 71.30 0.206 76.66 57.88 71.17 0.250
Mineru2-pipeline [14] - 75.51 0.209 76.55 70.90 79.11 0.225
PP-StructureV3 [10] - 86.73 0.073 85.79 81.68 89.48 0.073
General VLMs GPT-4o [7] - 75.02 0.217 79.70 67.07 76.09 0.148
InternVL3-76B [46] 76B 80.33 0.131 83.42 70.64 77.74 0.113
InternVL3.5-241B [47] 241B 82.67 0.142 87.23 75.00 81.28 0.125
Qwen2.5-VL-72B [24] 72B 87.02 0.094 88.27 82.15 86.22 0.102
Gemini-2.5 Pro [48] - 88.03 0.075 85.82 85.71 90.29 0.097
Specialized VLMs Dolphin [3] 322M 74.67 0.125 67.85 68.70 77.77 0.124
OCRFlux-3B [49] 3B 74.82 0.193 68.03 75.75 80.23 0.202
Mistral OCR [50] - 78.83 0.164 82.84 70.03 78.04 0.144
POINTS-Reader [4] 3B 80.98 0.134 79.20 77.13 81.66 0.145
olmOCR-7B [12] 7B 81.79 0.096 86.04 68.92 74.77 0.121
MinerU2-VLM [14] 0.9B 85.56 0.078 80.95 83.54 87.66 0.086
Nanonets-OCR-s [51] 3B 85.59 0.093 85.90 80.14 85.57 0.108
MonkeyOCR-pro-1.2B [1] 1.9B 86.96 0.084 85.02 84.24 89.02 0.130
MonkeyOCR-3B [1] 3.7B 87.13 0.075 87.45 81.39 85.92 0.129
dots.ocr [52] 3B 88.41 0.048 83.22 86.78 90.62 0.053
MonkeyOCR-pro-3B [1] 3.7B 88.85 0.075 87.25 86.78 90.63 0.128
MinerU2.5 [2] 1.2B 90.67 0.047 88.46 88.22 92.38 0.044
@@
Table 2: Comprehensive evaluation of document parsing on OmniDocBench v1.5. Results are reported by OmniDocBench [16] unless Ours.
@@
Pipeline Tools Docling-2.14.0 [11] 0.749 0.589 0.909 0.416 0.987 0.999 1 61.3 25.0 0.627 0.810 0.313 0.837
OpenParse-0.7.0 [53] 0.730 0.646 0.814 0.681 0.974 0.996 1 64.8 27.5 0.284 0.639 0.595 0.641
Unstructured-0.17.2 [54] 0.651 0.586 0.716 0.198 0.481 0.999 1 0 0.1 1 0.998 0.145 0.387
Pix2Text-1.1.2.3 [55] 0.424 0.320 0.528 0.138 0.356 0.276 0.611 73.6 66.2 0.584 0.645 0.281 0.499
Marker-1.7.1 [45] 0.397 0.296 0.497 0.085 0.293 0.374 0.688 67.6 54.0 0.609 0.678 0.116 0.329
Mathpix [56] 0.278 0.191 0.364 0.105 0.381 0.306 0.454 77.0 67.1 0.243 0.320 0.108 0.304
MinerU-pipeline [9] 0.203 0.162 0.244 0.072 0.111 0.313 0.581 77.4 79.5 0.166 0.150 0.097 0.136
PP-StructureV3 [10] 0.176 0.145 0.206 0.058 0.088 0.295 0.535 77.2 83.9 0.159 0.109 0.069 0.091
General VLMs InternVL2-76B [57] 0.442 0.440 0.443 0.353 0.290 0.543 0.701 63.0 60.2 0.547 0.555 0.317 0.228
GPT-4o [7] 0.316 0.233 0.399 0.144 0.409 0.425 0.606 72.0 62.9 0.234 0.329 0.128 0.251
InternVL3-78B [46] 0.257 0.218 0.296 0.117 0.210 0.380 0.533 69.0 73.9 0.279 0.282 0.095 0.161
Qwen2.5-VL-72B [24] 0.238 0.214 0.261 0.092 0.180 0.315 0.434 81.4 83.0 0.341 0.262 0.106 0.168
Gemini2.5-Pro [48] 0.180 0.148 0.212 0.055 0.168 0.356 0.439 85.8 86.4 0.130 0.119 0.049 0.121
Specialized VLMs Nougat [58] 0.713 0.452 0.973 0.365 0.998 0.488 0.941 39.9 0.0 0.572 1 0.382 0.954
SmolDocling-256M [13] 0.655 0.493 0.816 0.262 0.838 0.753 0.997 44.9 16.5 0.729 0.907 0.227 0.522
olmOCR-7B [12] 0.398 0.326 0.469 0.097 0.293 0.455 0.655 68.1 61.3 0.608 0.652 0.145 0.277
GOT [21] 0.349 0.287 0.411 0.189 0.315 0.360 0.528 53.2 47.2 0.459 0.520 0.141 0.280
OCRFlux-3B [49] 0.294 0.238 0.349 0.112 0.256 0.447 0.716 69.0 80.0 0.269 0.162 0.126 0.263
Nanonets-OCR-s [51] 0.289 0.283 0.295 0.134 0.231 0.518 0.546 76.8 79.4 0.343 0.201 0.135 0.200
Dolphin [3] 0.259 0.205 0.313 0.092 0.204 0.447 0.606 76.1 66.9 0.193 0.282 0.088 0.160
MinerU2-VLM [14] 0.186 0.133 0.238 0.045 0.115 0.273 0.506 82.1 83.4 0.150 0.209 0.066 0.122
MonkeyOCR-pro-1.2B [1] 0.184 0.146 0.221 0.068 0.118 0.272 0.452 81.3 85.5 0.149 0.134 0.093 0.179
MonkeyOCR-pro-3B [1] 0.172 0.138 0.206 0.067 0.107 0.246 0.421 81.5 87.5 0.139 0.111 0.100 0.185
dots.ocr [52] 0.143 0.125 0.160 0.032 0.066 0.329 0.416 88.6 89.0 0.099 0.092 0.040 0.067
MinerU2.5 [2] 0.143 0.111 0.174 0.050 0.074 0.258 0.473 88.3 89.2 0.089 0.083 0.045 0.068

### v4

@@ 1 Introduction
Documents serve as core information carriers, with their complexity and volume growing at an exponential rate, making document parsing an indispensable key technology. The primary goal of document parsing [li2025monkeyocr, niu2025mineru2, feng2025dolphin, liu2025points] is to enable deep structural and semantic understanding of a document’s layout. Specifically, it involves recognizing distinct text blocks and columns, distinguishing formulas, tables, charts, and images, determining the correct reading order, and detecting key elements (e.g., footnotes and image captions); these capabilities collectively lay a solid foundation for efficient information retrieval and data management. Furthermore, advanced document parsing enables large language models (LLMs) [ernie2025technicalreport, yang2025qwen3, achiam2023gpt], especially when combined with Retrieval-Augmented Generation (RAG) [lewis2020retrieval], to access high-quality knowledge and enhance their practical applications.
The inherent complexity of modern documents presents unique challenges: they often combine dense text, complex tables or chart, mathematical expressions, multiple languages and handwritten texts, with diverse layout structures. Recent research [li2025monkeyocr, wang2024mineru, cui2025paddleocr, Docling_Team_Docling, poznanski2025olmocr] in the field of document parsing primarily following two technological approaches. The first approach [wang2024mineru, cui2025paddleocr] employs pipeline methodologies based on specialized, modular expert models. Although these methods offer strong performance, they are increasingly hindered by integration complexity, cumulative error propagation, and inherent limitations when handling highly complex documents. Secondly, end-to-end approaches [poznanski2025olmocr, nassar2025smoldocling, MinerU2] leveraging multimodal models aim to simplify the workflow and enable joint optimization. However, these methods often struggle with correct text order and can even generate hallucinations when faced with lengthy or complex layouts, while also incurring substantial computational overhead for long sequence outputs, thereby restricting their practical deployment.
@@
Firstly, PaddleOCR-VL performs layout detection and reading order prediction to obtain the positional coordinates and reading order of elements (text blocks, tables, formulas, and charts). Compared to multimodal methods that rely on grounding and sequence output (e.g., MinerU2.5 [niu2025mineru2], Dolphin [feng2025dolphin]), our method offers faster inference speeds, lower training costs, and easier extensibility for new layout categories. Subsequently, the elements are segmented based on their positions and fed into PaddleOCR-VL-0.9B for recognition. This vision-language model is specifically designed for resource-efficient inference and excels at element recognition within document parsing. By integrating a NaViT-style [dehghani2023patch] dynamic high-resolution visual encoder with the lightweight ERNIE-4.5-0.3B [ernie2025technicalreport] language model, we have significantly enhanced the model’s dense text recognition capabilities and decoding efficiency.
@@
Comprehensive benchmarking on the public benchmarks, including OmniDocBench v1.0, v1.5 [ouyang2025omnidocbench] and olmOCR-Bench [poznanski2025olmocr], and in-house ones demonstrate that PaddleOCR-VL achieves SOTA performance in document parsing task, significantly outperforming existing pipeline-based solutions and exhibiting strong competitiveness against leading vision-language models (VLMs). Moreover, PaddleOCR-VL is optimized for efficiency, delivering substantially lower latency and higher throughput than competing approaches.
@@ 2.1.1 Layout Analysis
Specifically, we decouple the layout analysis process by introducing an independent model, PP-DocLayoutV2, dedicated solely to this task. PP-DocLayoutV2 consists of an object detection model (RT-DETR [zhao2024detrs]) for elements localization and classification, as well as a lightweight pointer network [hou2024relation] with six transformer layers to accurately predict the reading order of layout elements.
@@
Architecturally, PP‑DocLayoutV2 is composed of two sequentially connected networks, as shown in Figure 3. The first is an RT‑DETR-based [zhao2024detrs] detection model that performs layout element detection and classification. The detected bounding boxes and class labels are then passed to a subsequent pointer network, which is responsible for ordering these layout elements.
Specifically, we first apply per-class thresholds to select foreground proposals for the ordering network. The selected proposals are embedded using absolute 2D positional encodings and class label embeddings. Additionally, the encoder attention incorporates a geometric bias mechanism from Relation‑DETR [hou2024relation] to explicitly model pairwise geometric relationships among elements. The pairwise relation head linearly projects element representations into query and key vectors, then computes bilinear similarities to produce pairwise logits, resulting in an $N\times N$ matrix that represents the relative order between each pair of elements. Finally, a deterministic win‑accumulation decoding algorithm recovers a topologically consistent reading order for the detected layout elements.
In comparison to other specialized models, such as LayoutReader [wang2021layoutreader], our model achieves higher performance with fewer parameters by efficiently extending RT-DETR [zhao2024detrs] with a pointer network.
@@ 2.1.2 Element-level Recognition
We adopted an architectural style inspired by LLaVA [liu2023visual], integrating a pre-trained vision encoder with a dynamic resolution preprocessor, a randomly initialized 2-layer MLP projector, and a pre-trained large language model. Our architecture achieves a balance the scale of vision and language models to optimize performance in multi-elements recognition tasks.
Compared to earlier document parsing models based on fixed-resolution or tiling-based approaches [liu2025points, MinerU2, wei2024general], our approach utilizes native dynamic high-resolution preprocessing. For the vision encoder, we employed a NaViT-style [dehghani2023patch] encoder initialized from Keye-VL’s [team2025kwai] vision model, which support native-resolution inputs. This design enables the vision-language model to handle images of arbitrary resolution without distortion, yielding fewer hallucinations and stronger performance on text-intensive tasks.
The projector is a randomly initialized 2-layer MLP with GELU [hendrycks2016gaussian] activation, incorporating a merge size of 2 to efficiently bridge visual features from the encoder to the language model’s embedding space.
In auto-regressive language models, the entire sequence is generated by predicting one token at a time. This approach means that the size of the decoder is directly linked to the overall inference latency, so a smaller model will decode faster. With this in mind, we use the ERNIE-4.5-0.3B [ernie2025technicalreport] model, an open-source language model that balances a relatively small number of parameters with strong inference efficiency. In our implementation, we further enhance positional representation by incorporating a 3D-RoPE[bai2025qwen2].
The combination of NaViT [dehghani2023patch] with ERNIE-4.5-0.3B [ernie2025technicalreport] has led to significant performance improvements in documents parsing, achieving minimal memory usage and faster inference speed.
@@ 2.2.1 Layout Analysis
We employ the PP-DocLayoutV2 model to perform layout element localization, classification, and reading order prediction. PP-DocLayoutV2 extends RT-DETR [zhao2024detrs] by incorporating an additional pointer network [hou2024relation], which is responsible for predicting the reading order of detected elements. The training process adopts a two-stage strategy: we first train the core RT-DETR [zhao2024detrs] model for layout detection and classification. Afterward, we freeze its parameters and independently train the pointer network for reading order prediction.
For the first stage, we follow the training strategy of RT-DETR [zhao2024detrs]. Specifically, we initialize the model with PP-DocLayout_Plus-L [sun2025pp] pretrained weights and train it for 100 epochs on our self-constructed dataset comprising over 20,000 high-quality samples.
For the second stage, specifically, the model outputs a matrix representing the pairwise ordering relationships between any two elements, and the Generalized Cross Entropy Loss [zhang2018generalized] is computed with respect to the ground truth labels, as this loss function demonstrates increased robustness in scenarios where pre-annotated data are mixed into the dataset. We utilize a constant learning rate 2e-4 and the AdamW optimizer to train 200 epochs.
@@ 2.2.2 Element-level Recognition
As described in Section 2.1.2, PaddleOCR-VL-0.9B consists of three modules: a vision encoder, a projector, and a language model. We adopt a post-adaptation strategy using pre-trained models. Specifically, the vision model is initialized with Keye-VL’s weights, and the language model is initialized with ERNIE-4.5-0.3B’s weights. The model is trained based on the ERNIEKit [ERNIEkit] repository and the training methodology for our VLM is divided into two stages, as outlined in Table 1.
@@
    Table Recognition: The model learns to parse tabular structures within documents. This involves accurately extracting cell contents, identifying rows and columns, and recognize the logical relationships between different table elements, ultimately generating structured representations based on OTSL [lysak2023optimized] format.
@@ 3.1 Data Curation
    Open Source Dataset: As the foundation of our dataset, we systematically aggregated and curated a wide array of established public datasets. For textual content, we sourced data from the canonical dataset CASIA-HWDB [liu2011casia]. Our mathematical expression data is derived from UniMER-1M [unimernet] and MathWriting [MathWriting]. To ensure comprehensive coverage of data visualizations, we incorporated a rich spectrum of chart and graph datasets, including ChartQA [chartqa], PlotQA [plotqa], Chart2Text [chart2text], DVQA [dvqa], Unichart [masry2023unichart], Beagle [beagle], ChartINFO [chartinfo], visText [tang2023vistext], and ExcelChart [excelchart]. Each of these sources underwent an initial filtering and cleaning protocol to rectify or discard noisy and low-quality annotations.
@@ 3.2 Automatic Data Annotation
After acquiring the raw data, we utilize an automatic data annotations process for large-scale labeling. Initially, we employ the expert model, PP-StructureV3, to conduct preliminary processing on the data, generating pseudo labels that may contain some inaccuracies. Subsequently, through prompt engineering, we create prompts that include the original images and their associated pseudo labels, which are then submitted to more advanced multimodal large language models, ERNIE-4.5-VL [ernie2025technicalreport] and Qwen2.5VL [bai2025qwen2]. These sophisticated models refine and enhance the initial results by analyzing the image content, resulting in improved labels. Finally, to ensure the quality of the labels, the system performs a hallucination filtering step, which eliminates any potentially incorrect content generated by the large models, thereby producing reliable and high-quality labels.
@@ 3.3 Hard Cases Mining
To overcome performance bottlenecks in specific complex scenarios, we propose a hard case mining process for targeted performance improvement. We firstly develop a eval engine for various types. We created substantial evaluation data with precisely labeled data obtained through manual annotation. Theses evaluation datasets are categorized into several types: text data includes 23 categories such as Chinese, English, printed, handwritten, Japanese, Latin, and emojis; table data includes 20 categories such as limited tables, unlimited tables, handwritten tables, checklists, invoices, and rotated tables; formula data includes 4 categories such as Chinese and English formulas, handwritten and printed, simple, and complex; chart data includes 11 categories such as Chinese and English charts, line charts, and bar charts, sourced from diverse origins to cover different document. By inference on this evaluation set and using corresponding professional metrics (e.g., EditDist for Text, TEDS [teds] for Tables, RMS-F1 [deplot] for Charts, and BLEU [bleu_score] for Formulas), we can accurately identify hard cases where the model performs poorly. Finally, for these identified weaknesses, the system utilizes a rich set of resources (such as Font Library, CSS Library, Corpus) and rendering tools (like XeLaTeX and web browsers) to synthetically generate a large volume of new, high-quality hard cases.
@@ 4.1 Page-level Evaluation
To comprehensively evaluate the document parsing capabilities, we conducted extensive experiments on the OmniDocBench v1.5 [niu2025mineru2] benchmark. It is an expansion of version v1.0, adding 374 new documents for a total of 1,355 document pages. It features a more balanced distribution of data in both Chinese and English, as well as a richer inclusion of formulas and other elements. The evaluation method has been updated, with formulas assessed using the CDM method. The overall metric is a weighted combination of the metrics for text, formulas, and tables.
Table 2 demonstrate that PaddleOCR-VL achieves SOTA performance, outperforming existing pipeline tools, general VLMs, and other specialized document parsing models across all key metrics. Specifically, our model achieves a top-ranking overall score of 92.86, surpassing the next best model, MinerU2.5-1.2B (90.67). Moreover, our model establishes new SOTA results in the sub-tasks, including the lowest Text-Edit distance [lcvenshtcin1966binary] of 0.035, the highest Formula-CDM score of 91.22, the leading scores of 90.89 and 94.76 in Table-TEDS and Table-TEDS-S, and the best readering ordering scores of 0.043, respectively. These results underscore its superior accuracy in text recognition, formula recognition, and complex table structure analysis.
@@
Pipeline Tools Marker-1.8.2 [vik2024marker] - 71.30 0.206 76.66 57.88 71.17 0.250
Mineru2-pipeline [MinerU2] - 75.51 0.209 76.55 70.90 79.11 0.225
PP-StructureV3 [cui2025paddleocr] - 86.73 0.073 85.79 81.68 89.48 0.073
General VLMs GPT-4o [achiam2023gpt] - 75.02 0.217 79.70 67.07 76.09 0.148
InternVL3-76B [zhu2025internvl3] 76B 80.33 0.131 83.42 70.64 77.74 0.113
InternVL3.5-241B [wang2025internvl35] 241B 82.67 0.142 87.23 75.00 81.28 0.125
Qwen2.5-VL-72B [bai2025qwen2] 72B 87.02 0.094 88.27 82.15 86.22 0.102
Gemini-2.5 Pro [gemini25] - 88.03 0.075 85.82 85.71 90.29 0.097
Specialized VLMs Dolphin [feng2025dolphin] 322M 74.67 0.125 67.85 68.70 77.77 0.124
OCRFlux-3B [OCRFlux2025] 3B 74.82 0.193 68.03 75.75 80.23 0.202
Mistral OCR [mistral] - 78.83 0.164 82.84 70.03 78.04 0.144
POINTS-Reader [liu2025points] 3B 80.98 0.134 79.20 77.13 81.66 0.145
olmOCR-7B [poznanski2025olmocr] 7B 81.79 0.096 86.04 68.92 74.77 0.121
MinerU2-VLM [MinerU2] 0.9B 85.56 0.078 80.95 83.54 87.66 0.086
Nanonets-OCR-s [Nanonets-OCR-S] 3B 85.59 0.093 85.90 80.14 85.57 0.108
MonkeyOCR-pro-1.2B [li2025monkeyocr] 1.9B 86.96 0.084 85.02 84.24 89.02 0.130
MonkeyOCR-3B [li2025monkeyocr] 3.7B 87.13 0.075 87.45 81.39 85.92 0.129
dots.ocr [dotsocr] 3B 88.41 0.048 83.22 86.78 90.62 0.053
MonkeyOCR-pro-3B [li2025monkeyocr] 3.7B 88.85 0.075 87.25 86.78 90.63 0.128
MinerU2.5 [niu2025mineru2] 1.2B 90.67 0.047 88.46 88.22 92.38 0.044
@@
Table 2: Comprehensive evaluation of document parsing on OmniDocBench v1.5. Results are reported by OmniDocBench [ouyang2025omnidocbench] unless Ours.
@@
Pipeline Tools Docling-2.14.0 [Docling_Team_Docling] 0.749 0.589 0.909 0.416 0.987 0.999 1 61.3 25.0 0.627 0.810 0.313 0.837
OpenParse-0.7.0 [open-parse] 0.730 0.646 0.814 0.681 0.974 0.996 1 64.8 27.5 0.284 0.639 0.595 0.641
Unstructured-0.17.2 [unstructured] 0.651 0.586 0.716 0.198 0.481 0.999 1 0 0.1 1 0.998 0.145 0.387
Pix2Text-1.1.2.3 [Pix2Text] 0.424 0.320 0.528 0.138 0.356 0.276 0.611 73.6 66.2 0.584 0.645 0.281 0.499
Marker-1.7.1 [vik2024marker] 0.397 0.296 0.497 0.085 0.293 0.374 0.688 67.6 54.0 0.609 0.678 0.116 0.329
Mathpix [mathpix] 0.278 0.191 0.364 0.105 0.381 0.306 0.454 77.0 67.1 0.243 0.320 0.108 0.304
MinerU-pipeline [wang2024mineru] 0.203 0.162 0.244 0.072 0.111 0.313 0.581 77.4 79.5 0.166 0.150 0.097 0.136
PP-StructureV3 [cui2025paddleocr] 0.176 0.145 0.206 0.058 0.088 0.295 0.535 77.2 83.9 0.159 0.109 0.069 0.091
General VLMs InternVL2-76B [InternVL] 0.442 0.440 0.443 0.353 0.290 0.543 0.701 63.0 60.2 0.547 0.555 0.317 0.228
GPT-4o [achiam2023gpt] 0.316 0.233 0.399 0.144 0.409 0.425 0.606 72.0 62.9 0.234 0.329 0.128 0.251
InternVL3-78B [zhu2025internvl3] 0.257 0.218 0.296 0.117 0.210 0.380 0.533 69.0 73.9 0.279 0.282 0.095 0.161
Qwen2.5-VL-72B [bai2025qwen2] 0.238 0.214 0.261 0.092 0.180 0.315 0.434 81.4 83.0 0.341 0.262 0.106 0.168
Gemini2.5-Pro [gemini25] 0.180 0.148 0.212 0.055 0.168 0.356 0.439 85.8 86.4 0.130 0.119 0.049 0.121
Specialized VLMs Nougat [blecher2023nougat] 0.713 0.452 0.973 0.365 0.998 0.488 0.941 39.9 0.0 0.572 1 0.382 0.954
SmolDocling-256M [nassar2025smoldocling] 0.655 0.493 0.816 0.262 0.838 0.753 0.997 44.9 16.5 0.729 0.907 0.227 0.522
olmOCR-7B [poznanski2025olmocr] 0.398 0.326 0.469 0.097 0.293 0.455 0.655 68.1 61.3 0.608 0.652 0.145 0.277
GOT [wei2024general] 0.349 0.287 0.411 0.189 0.315 0.360 0.528 53.2 47.2 0.459 0.520 0.141 0.280
OCRFlux-3B [OCRFlux2025] 0.294 0.238 0.349 0.112 0.256 0.447 0.716 69.0 80.0 0.269 0.162 0.126 0.263
Nanonets-OCR-s [Nanonets-OCR-S] 0.289 0.283 0.295 0.134 0.231 0.518 0.546 76.8 79.4 0.343 0.201 0.135 0.200
Dolphin [feng2025dolphin] 0.259 0.205 0.313 0.092 0.204 0.447 0.606 76.1 66.9 0.193 0.282 0.088 0.160
MinerU2-VLM [MinerU2] 0.186 0.133 0.238 0.045 0.115 0.273 0.506 82.1 83.4 0.150 0.209 0.066 0.122
MonkeyOCR-pro-1.2B [li2025monkeyocr] 0.184 0.146 0.221 0.068 0.118 0.272 0.452 81.3 85.5 0.149 0.134 0.093 0.179
MonkeyOCR-pro-3B [li2025monkeyocr] 0.172 0.138 0.206 0.067 0.107 0.246 0.421 81.5 87.5 0.139 0.111 0.100 0.185
dots.ocr [dotsocr] 0.143 0.125 0.160 0.032 0.066 0.329 0.416 88.6 89.0 0.099 0.092 0.040 0.067
MinerU2.5 [niu2025mineru2] 0.143 0.111 0.174 0.050 0.074 0.258 0.473 88.3 89.2 0.089 0.083 0.045 0.068